OpenStudy (anonymous):

tanx cosx-cos^2x=sin^2x

OpenStudy (e.mccormick):

Som a proof?

OpenStudy (e.mccormick):

\[\tan x \cos x-\cos^2x=\sin^2x \implies \\ \frac{\sin x}{\cos x} \cos x-\cos^2x=\sin^2x\implies \] That is the typical start of these, which you can see gets rid of the cosx, but what would you do next? I know what I would try, but where would you head?

OpenStudy (e.mccormick):

And could you check to make sure you wrote the problem correctly? Just wondering about something.

OpenStudy (anonymous):

ok